OpenStudy (anonymous):

a cell of the bacterium escherichia coli in a nutrient-broth divided into two cells every 20 min. Suppose that there are initially 500 cells . Find the number of cells after t hours. How many cells will there e after 8 hours?

OpenStudy (mertsj):

Oh let me think...isn't that equation \[A=A _{0}e ^{kt}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but I use a instead of e since the equation with e is for "continuously" questions

OpenStudy (mertsj):

So first we need to find k. And we know that if the original amount is 1, in 1/3 hour, there will be 2

OpenStudy (mertsj):

Could you post your equation?

OpenStudy (anonymous):

A=A0 a ^ik

OpenStudy (mertsj):

What is a?

OpenStudy (mertsj):

What does it represent?

OpenStudy (mertsj):

Would it be 2 in this example?

OpenStudy (mertsj):

And i is the time period?

OpenStudy (anonymous):

I think it represents if the question is a half life question or doubling time, so if its half life, then it would be 1/2, so i think in this it would be 2 since the bacteria doubles

OpenStudy (mertsj):

\[2=1(2)^{\frac{1}{3}k}\]

OpenStudy (anonymous):

t is the time, its like the x-value . and i is the doubling time

OpenStudy (anonymous):

ok, so why would I not use the 500 initial amount yet, because when I did it i used 500 instead of 1 to find the equation

OpenStudy (mertsj):

\[\ln 2=\frac{1}{3}k \ln 2\]

OpenStudy (mertsj):

k=3

OpenStudy (mertsj):

\[A=500(2)^{8(3)}\] \[A=8,388,608,000\]

OpenStudy (anonymous):

every time i sub 8 into the equation i don't get the answer you have, which is the right answer

OpenStudy (mertsj):

Did you get 3 for k?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Did you do 2^24 first?

OpenStudy (mertsj):

And then times 500?

OpenStudy (anonymous):

no. I did the same questions in another math class and I was able to put the whole equation into my cal and it would give me the right answer, but now i get the wrong one

OpenStudy (mertsj):

I don't know about your calculator but I know you do powers first. Do do 2^24 and take that answer times 500

OpenStudy (anonymous):

ok. Could you answer another question?

OpenStudy (mertsj):

Possibly

OpenStudy (anonymous):

The count in a bacteria culture was 5000 after 15 min and 40 000 after 1 hour. What was the initial size of the culture. The two equations I came up with are 5000=A(2)^1/4t and 40000=A(2)^t

OpenStudy (mertsj):

\[40000=5000(2)^{\frac{3}{4}k}\]

OpenStudy (mertsj):

Find k

OpenStudy (anonymous):

how did you get your equation, because my teacher said to either make the two equation equal each other and solve for k, or use substitution and when I did both i got wrong answers

OpenStudy (mertsj):

If there were 5000 after 15 minutes, there were 10,000 after 30 minutes, 20,000 after 45 minutes, and 40,000 after 60 minutes. So if we take 40,000 for the final amount and 5000 for the beginning amount, and 3/4 hour for the time, we have the equation I posted.

OpenStudy (mertsj):

The problem I see with your equations is that you don't know what k is.

OpenStudy (mertsj):

If you wanted two equations you could write: \[5000=A _{0}(2)^{\frac{1}{4}k}\] and \[40000=A _{0}(2)^{1k}\]

OpenStudy (anonymous):

how would I solve it if I use substitution, because when I was doing it I was dividing out the 2

OpenStudy (mertsj):

I would divide the second equation by the first:

OpenStudy (mertsj):

\[\frac{40,000}{5,000}=\frac{A _{0}(2)^{1k}}{A _{0}(2)^{\frac{1}{4}k}}\]

OpenStudy (mertsj):

\[8=2^{\frac{3}{4}k}\]

OpenStudy (mertsj):

Then solve for k

OpenStudy (anonymous):

would the 2 in the two equations not cancel out?

OpenStudy (mertsj):

After you have k, you can find A_0 by substitution

OpenStudy (mertsj):

No. When you do this problem: \[\frac{x^8}{x^3}\] do the x's cancel out?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

Then why would the 2's cancel out?

OpenStudy (anonymous):

i don't know. Sorry, i'm thinking of the 2 without the exponent. I'm sorry. I understand why now

OpenStudy (anonymous):

could you stay around if i have another question

OpenStudy (mertsj):

Did you get this one finished?

OpenStudy (anonymous):

almost, it has sub-parts . working on it now

OpenStudy (anonymous):

one of the sub-parts asks to find the population after t hours, the answer is 2500(16)^t

OpenStudy (anonymous):

i have 2500(2)^4t

OpenStudy (anonymous):

would they both be right?

OpenStudy (mertsj):

I don't understand why you have 16 for the base in the one equation.

OpenStudy (mertsj):

What did you get for k?

OpenStudy (anonymous):

the answer at the back of my book has 16 as the base, I got 4 for k

OpenStudy (mertsj):

And I take it 2500 is the beginning amount. The reason they have 16 for the base is that they incorporated the constant k into the base:

OpenStudy (mertsj):

\[A=2500(2)^{4t}=2500(2^4)^t=2500(16)^t\]

OpenStudy (anonymous):

would my answer of 2500 (2)4t be right?

OpenStudy (mertsj):

Yes. They are the same thing.

OpenStudy (mertsj):

Don't you understand what I just posted?

OpenStudy (anonymous):

I did, I just need to make sure I'm doing it right because I suck at word problems. I'm sorry for being so annoying..

OpenStudy (mertsj):

It's ok.

OpenStudy (anonymous):

one of my questions says that the world population doubles every 35 a. What does it mean by a, I tried to look it up but I didn't find anything. Does it mean years